OpenStudy (anonymous):

The length of a rectandle is 8 inches greater than it's width. Find the width of the rectangle if its area is 105 square inches.

mathslover (mathslover):

Ok ! first of all remember that area of a rectangle : length * width given length = 8 inches + width area = 105 inch^2 \[105 inch^2 = (8 inch + width)width = 8inch * width + width^2\]

OpenStudy (anonymous):

Dude.. I'm freaking lost. You completely just confused me.

OpenStudy (anonymous):

Dude.. I'm freaking lost. You completely just confused me.

mathslover (mathslover):

let width be x : \[\huge{8x+x^2=105^2}\]

mathslover (mathslover):

is this equation ok to you ?

OpenStudy (anonymous):

 area is x(x+8) but area is 105 x(x+8)=105